# Quantum Geometry: Revisiting electronic scales in quantum matter

### v1

@@ V Discussion and Outlook
Let us conclude with the question: is it necessary to include quantum geometry in the standard toolbox? As physicists, we often look for the simplest description that abstracts away irrelevant degrees of freedom. Quantum geometry quantifies this process by allowing us to define energy and length scales with origin in orbital mixing. In modeling any condensed matter system, we strive to find a middle ground between complexity and simplicity, a compromise that dictates the level of description we adopt. Indeed, there is no ambiguity in defining the many-body Hamiltonian at the UV scale - the kinetic energy depends on the momentum operator, and interactions depend on the density operator. It is only when the theory is projected onto a subset of states that quantum geometric features emerge, often leading to low-energy theories that defy naive semiclassical intuition, as seen in flat-band superconductors and fractionalized quantum Hall states [143]
Quantum geometry is thus a natural and essential byproduct of our reliance on effective theories. Tying back to the notion of dipole fluctuations, this perspective can be succinctly stated as such: If the ground state of a material has strong orbital mixing, it cannot be described simultaneously on the basis of a single local orbital and in flat space. We can either choose a multi-band picture in flat space or a single-band picture in a curved geometry.

### v2

@@ Introduction
Let us conclude with the question: is it necessary to include quantum geometry in the standard toolbox? As physicists, we often look for the simplest description that abstracts away irrelevant degrees of freedom. Quantum geometry quantifies this process by allowing us to define energy and length scales with origins in orbital mixing. In modeling any condensed matter system, we strive to find a middle ground between complexity and simplicity, a compromise that dictates the level of description we adopt. Indeed, there is no ambiguity in defining the many-body Hamiltonian at the UV scale - the kinetic energy depends on the momentum operator, and interactions depend on the density operator. It is only when the theory is projected onto a subset of states that quantum geometric features emerge, often leading to low-energy theories that defy naive semiclassical intuition, as seen in flat-band superconductors and fractionalized quantum Hall states [157]. Quantum geometry is thus a natural and essential byproduct of our reliance on effective theories. Tying back to the notion of dipole fluctuations, this perspective can be succinctly summarized as follows: If the ground state of a material has strong orbital mixing, it cannot be described simultaneously on the basis of a single local orbital and in flat space. We can either choose a multi-band picture in flat space or a single-band picture in a curved geometry.
Box 1: A case study of transition metal dichalcogenides
Transition metal dichalcogenide semiconductors exemplify many of the central themes of this review. They share the chemical formula $\rm MX_{2}$, where $\rm M$ is a transition metal ($\rm Mo$, $\rm W$) and $\rm X$ is a chalcogen ($\rm S$, $\rm Se$, $\rm Te$). Each monolayer consists of a hexagonal lattice of transition metal atoms sandwiched between two chalcogen planes, forming a quasi-2D crystal with strong in-plane covalent bonds and weak interlayer van der Waals interactions. For broad overviews of their chemistry, band structure, and optical properties, see Refs. [158, 159]. Like graphene, whose physical properties are well understood to stem from the geometry of its Dirac points, these materials show a remarkable zoo of exotic correlated phases and responses which can be attributed to the nontrivial structure of their wavefunctions.
Unlike semiconductors such as GaAs where the effective mass approximation applies well, minimal tight-binding models of transition metal dichalcogenides cannot be reduced to a single-orbital. Rather, the conduction and valence bands derive from the bonding of three hybrid orbitals across neighboring unit cells, localized mostly in the empty space between metal and chalcogen [160]. This originates from an effective lattice frustration and results in a substantial flattening of the valence band. In these semiconductors, dipole fluctuations are large and extended beyond one unit cell. This topological obstruction [161] produces very large valley-contrasting Berry curvature [162, 163], spin–orbit splitting [164, 165], and a nontrivial band geometry[166].
Because of this intrinsic geometry, both Landau levels [167, 168] and excitons [169] have long been observed to deviate from the effective mass approximation, with some of these deviations attributed to quantum geometric contributions [170, 171]. In the optical domain, this same quantum geometry underlies the giant nonlinear responses [172], including unusually strong second-harmonic generation [173]. All these effects can, retrospectively, be attributed to abnormally large dipole fluctuations, due to their multi-orbital character.
Twisting or stacking these materials into moiré heterostructures amplifies these effects by producing flat bands in which geometric contributions dominate the interactions. Experimentally, such systems host a remarkably wide range of correlated phenomena, including superconductivity [174], fractional quantum Hall phases [126, 127, 175, 176], quantum anomalous Hall effects [177], generalized Wigner crystals [178, 179], and other symmetry-broken insulating states [180, 181, 182, 183, 184]. On the theoretical side, the role of quantum geometry in these moiré bands has been elucidated through studies of form factors [185, 186, 187, 188, 189, 190] and chiral limits [191]. For a broader overview of this rapidly growing field we refer readers to Ref. [4].
Acknowledgments
R.Q. and N.V. acknowledge support from the U.S. Department of Energy through the Center for Programmable Quantum Materials (DE-SC0019443). R.Q. also acknowledges support from the National Science Foundation CAREER program (DMR-2340394) and the Alfred P. Sloan Foundation (FG-2025-24714). PJWM acknowledges support by the European Research Council (ERC) under the European Union’s Horizon research and innovation programme (XBEND, grant agreement no. 101080740). T.H. acknowledges financial support by the European Research Council (ERC) under grant QuantumCUSP (Grant Agreement No. 101077020).
Author contributions
All authors contributed to the conceptualization, writing, and editing of the manuscript.